Discuss with Askiitians Tutors

I have recently purchased pendrive video lectures of iit jee from askiitians.My 11 and 12th concepts are not clear.I want a time table to crack iit jee 2017 in 3 months,please help me.I am not going to any school or coaching

Ram Singh
9 Years agoGrade

1 Answer

Umakant biswal
9 Years ago
@ mihir 
so, u are a dropper .. ryt ?? if so, i am giving u time table below , please stick to it constantly for better ranks . 
5;30-6;30 = YOGA 
6;30 – 7;30 – DAILY WORK ( LIKE BRUSHING , BATHING ) 
7;30 – 7;45 – MORNING BREAKFAST 
8;00 – 10;00 – MATHS ( As this time is suitable for maths ) 
10;00- 10;15 – break 
10;16- 11;10 – previous year paper practice of whatever u studied in maths 
11;10- 11;20 – break 
11;20 – 1;15 – physics ( theory 1 hr , numerical 1 hr ) 
if u feel hungry iun between then u can use break time to eat something . 
1;15- 1;45= mid day meals 
1;45 – 2;10 – watch tv if u intrested in it . 
2;10 – 4;15 = rest 
4;15 – 4;40 – play outdoor games 
4;40 – 5;00 – take some break fast 
5;00 – 7;00 = chemistry( physical 2 days , inorgaNIC 2 days , and organic 3 days per week ) 
7;00 – 7;’15 – break 
7;15 = 9;15= revision of whatever u have studied whole day ( 1 hr revision , 1 hr prev year practice ) 
then after 9 ; 15 spent some time to read magazines for iitjee like chemistry spectrum , physics spectrum etc .. 
ALL THE BEST ..